What is the amount of heat generated during the freezing of m=10kg of water taken at temperature T=0°c.The heat of ice melting is 3.3×10⁴Jkg-¹
1
Expert's answer
2020-08-10T19:54:26-0400

The amount of heat generated during the freezing of water at temperature T=0°c can be calculated with the following formula:


"Q = \lambda m"

where "m = 10 kg" is the mass of water and "\lambda = 3.3\times 10^4 J/kg" is the heat of ice melting. Hence, obtain:


"Q = \lambda m = 3.3\times 10^4\cdot 10 = 3.3\times 10^5J = 0.33\times 10^6J = 0.33MJ"

Answer. 3.3×10^5 J or 0.33 MJ.